OpenStudy (anonymous):

The frame for a rectangular box is to be constructed from 44 feet of a thin metal rod. The metal rod is to be cut into pieces such that the bottom of the box is a square with edge length x. Express the volume, V, of the box as a function of x.

OpenStudy (cwrw238):

V = x^2 * h also total length of frame = 44 = 4x + 4x + 4h 44 = 8x + 4h express h in terms of x using above equation and substitute for h in the formula for V

OpenStudy (anonymous):

thank you))

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

can you continue from the above?

OpenStudy (anonymous):

I guess...but could you please finish it?

OpenStudy (campbell_st):

the square length is x  there is 8x in the frame height = (44 - 8x)/4 4 edges of the same height height = 11 - 2x the volume is \[V = x^2 \times (11 - 2x) = 11x^2 - 2x^3\]